OpenStudy (anonymous):

Rewrite 0.ababababab.... as a repeating fraction

OpenStudy (anonymous):

It tells me that I got it wrong, but I will show my work here

OpenStudy (rational):

try (10a+b)/99

OpenStudy (anonymous):

\[\Large \color{red}{x}=0.ababab...\] \[\Large \color{blue}{100x}=ab.ababab...\] \[\Large \color{blue}{100x}=ab.ababab...\] - \[\Large \color{red}{x}=00.ababab...\] \[\Large \color{green}{99x}=ab\] \[\Huge x=\frac{ ab }{ 99 }\]

OpenStudy (rational):

You're rihgt, just notice that ab is not a*b "ab" is just the positional notation in decimal number system : \[x = \dfrac{(ab)_{10}}{99} = \dfrac{10a+b}{99}\]

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

Oh, since a is in the 10s place, its actually 10a and b is in the ones place?

OpenStudy (anonymous):

for example \[0.131313...=\frac{13}{99}\]

OpenStudy (anonymous):

I understand now.